convert the following equations into logrimithic form
6=2x
-9=6y
7=3y
x=5y

Answer by stanbon(75887)

convert the following equations into logrimithic form
6=2^x
log(base 2) 6 = x
--------------------------

-9=6^y
log(base 6) -9 = y
Comment: That is meaningless as no power of 6 can give you a negative number.
--------------------------

7=3^y
log(base 3) 7 = y
--------------------------

x=5^y
log(base 5) x = y
=============================
Cheers,
Stan H.